OpenStudy (anonymous):

A sinusoidal function passes through (0,2), has a max point at (20 degrees,5), and a min point at (60 degrees, -1). How do you determine a possible sine and cosine function once you determine the amplitude, period and range of the function

OpenStudy (anonymous):

Ok so the amplitude is the mid point. Its the Highest point on the y axis - The lowest point on the y axis and then divided by 2 So the Highest point is 5 and the lowest point is -1 \[ \text{ Amplitude } = \frac{5-(-1)}{2}= \frac{6}{2}=3\]

OpenStudy (anonymous):

Now the maximum point to the minimum point on the graph is only half the period/cycle